# On the Ability of Deep Networks to Learn Symmetries from Data: A Neural Kernel Theory

### v1

@@ 3 Learning from partially observed symmetries
Figure 1: Common deep network architectures fail to extrapolate symmetries from a partially observed version of rotated-MNIST. A: A conceptual sketch of the learning task. For samples in the leave-out class (digit “5” in this example), the upright pose is not included in the training set (white circles). For samples in the remaining classes, all poses are included in the training set (black circles). A model that can generalize rotations should classify the missing upright samples correctly. B: The accuracies of the three models tested respectively on the training, in-test set (normal i.i.d. test set) and out-test set (i.e., left-out pose of the leave-out class). Chance levels (10%) are also reported as a baseline. Error shades represent 95% confidence intervals computed over all tested leave-out classes (n = 10). C: Out-test accuracy for the three models as a function of the number of angles in the rotation orbit. Error shades as in B.
3.1 Illustration of a symmetry learning problem
@@
To illustrate the problem of symmetry learning, we train a range of networks with different architectures (MLP, ConvNet, ViT-S (Dosovitskiy,, 2020)) on partially observed versions of rotated-MNIST (details of the architectures in App. B).
@@
We start by constructing the rotated-MNIST dataset: for all samples in MNIST, we generate $N$ rotated samples, where the rotation angles are multiples of $2\pi/N$. We choose a digit class, which we refer to as the leave-out class (Fig. 1A). We define the “training” and the “in-test” splits as a 90/10 random split of the whole rotated MNIST dataset, except the upright samples of the leave-out class. These samples, instead, constitute what we call the “out-test” split. Note that predicting the labels of the out-test split is an out-of-distribution (OOD) generalization task. In this work, we study whether networks are able to perform this OOD task in virtue of the symmetric nature of the dataset.
@@
We next show that increasing the number of angles leads to better out-test performance across models (Fig. 1C). Successful generalization is achieved when the number of angles sampled is sufficiently large.
In the following, we build a theoretical framework to understand and predict the generalization abilities of deep networks on problems of this type, involving learning from partially observed symmetries. Fully developed, our theory gives a precise account of the generalization behavior of normally trained networks on rotated-MNIST observed here (Section 3.4 and Fig. 6).
@@ 3.2 A simple case study: Gaussian kernel regression on a circular dataset
As a first step toward our goal, we study a simple scenario which captures the relevant phenomenology for understanding the behavior of deep networks on datasets with symmetries.
Circular dataset.
Consider the following dataset of $2N$ points:
$\mathcal{D}=\{(g^{i}.x^{A},+1)\}_{i=0}^{N-1}\cup\{(g^{i}.x^{B},-1)\}_{i=0}^{N-1}\subset\mathbb{R}^{n}\times\mathbb{R},$ (1)
@@
where $g$ is the representation of the generator of the cyclic group of order $N$, acting on $\mathbb{R}^{n}$ by usual matrix multiplication. Such dataset is composed of two orbits of the same group, obtained for two different “seed” samples, $x^{A}$ and $x^{B}$, which are labeled +1 and -1 respectively. In the following, we will use this type of dataset as a blueprint, and specialize it to different case studies. For reasons that will become apparent soon, we order the dataset so that points labeled +1 are interleaved with those labeled -1, while preserving individual orbit ordering. For notational convenience, we denote $x_{i}^{\bullet}=g^{i}.x^{\bullet}$. The ordered dataset can be written:
$\mathcal{D}_{o}$ =$=\{(g^{0}.x^{A},+1),(g^{0}.x^{B},-1),(g^{1}.x^{A},+1),\cdots(g^{N-1}.x^{B},-1)\}$ (2)
=$=\{(x_{0}^{A},+1),(x_{0}^{B},-1),(x_{1}^{A},+1),\cdots(x_{N-1}^{B},-1)\}.$ (3)
In this section, we begin with a simple, low dimensional instantiation of the dataset $\mathcal{D}_{o}$, where the two orbits are discretely sampled and interleaved circles of radius 1, separated by a distance $\Delta$ (see Fig. 2A for an illustration). The following definitions of $g$, $x^{A}$ and $x^{B}$ are thus only valid in the context of this section (Section 3.2). We consider the following $\mathbb{R}^{3}$ representation of the generator $g$:
$g=\begin{bmatrix}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{bmatrix},$ (4)
where we define $\theta=2\pi/N$. We consider also the seed points
$x^{A}=\frac{\Delta}{2}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}+\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},\quad x^{B}=-\frac{\Delta}{2}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}+\begin{bmatrix}0\\ \cos(\theta/2)\\ \sin(\theta/2)\end{bmatrix},$ (5)
@@
over which we make $g$ act by usual matrix multiplication. Successive applications of $g$ onto the seed points generate the dataset. The second seed point $x^{B}$ is chosen such that the two orbits are geometrically interleaved.
@@
Figure 2: Case study: Analysis of the generalization behavior of the Gaussian kernel on a circular-symmetric dataset. A: A circular dataset is made of two sets of interleaved points in $\mathbb{R}^{3}$ belonging to two classes (denoted in red and blue). One point (in black) is left out during training, and we regress on it. B: The pairwise distance matrix between the points is circulant (above), leading to a circulant kernel matrix (below) obtained by applying the Gaussian (RBF) kernel function to it elementwise. We select the $N$-th row from it, highlighted by the red rectangle; due to the circularity of the matrix, this comes without loss of information. C: A plot of the selected row of the kernel matrix. The values of the kernel alternate between two limiting curves, representing respectively the “same label” and the “opposite label” kernel values. This alternation, due to the separation between classes, is akin to the presence of a high frequency component in the kernel function itself, were it computed over the classes perfectly interleaved ($\Delta$ = 0). D: The prediction error of a Gaussian kernel on a missing point of a circular-symmetric dataset is a simple function of its spectrum (Eq. 7). The reciprocal (inverse) of the positive half of the DFT of the selected row of the kernel matrix is shown. The grey area corresponds to the numerator of the spectral error, while the red area corresponds to the denominator. The ratio of the two corresponds to the prediction error incurred by GP regression using the chosen kernel. The presence of the aforementioned high frequency component in the kernel is manifested in the low value of $\lambda_{N}^{-1}$, leading to small error (i.e., increased class separability).
@@
Figure 3: Further geometric interpretations of the Gaussian kernel generalization behavior on a circular-symmetric dataset. A: We progressively increase the separation $\Delta$ between two circular classes of points (respectively in blue and red). The leave-out point used for testing is in black. B: The prediction error (in black) of the RBF kernel on the leave-out point decreases as a function of $\Delta$, as predicted by our formula (in red) given in Eq. 7 relating generalization error to spectral quantities of the kernel. C: Inverse spectra of the kernel matrix for different $\Delta$. The grey (respectively, red) area corresponds to the numerator (resp., denominator) of Eq. 7. As $\Delta$ increases, the ratio between the grey and the red areas progressively decreases as a consequence of the last frequencies getting larger. D, E, F: Same as A,B,C where instead of increasing $\Delta$ we increase the number of points forming each class. We see that the ratio between grey and red areas (i.e., the spectral error) progressively decreases with number of points, as a consequence of the middle inverse frequencies progressively diverging as point density increases.
Gaussian kernel regression on a circular dataset.
@@
We consider the following regression problem: remove any one point from the dataset $\mathcal{D}_{o}$, and find a predictor to estimate the value of $y$ at the missing point. We solve this regression problem using Gaussian process (GP) regression. We select the Gaussian or Radial Basis Function (RBF) kernel,
$k_{RBF}(x_{i},x_{j})=\exp(-L^{2}\lVert x_{i}-x_{j}\rVert_{2}^{2}),$ (6)
where $L$ denotes the kernel’s length scale. We use this kernel function to compute the kernel matrix (or Gram matrix) of our dataset, $K_{train}\in\mathbb{R}^{(2N-1)\times(2N-1)}$. The dataset is constructed and ordered in such a way that the pairwise distance matrix computed over points of the dataset is circulant (Fig. 2B). The reader can convince themselves of this by noticing that the pairwise distances between points $x_{i}$ and $x_{j}$ in the circular orbits are only function of their absolute index difference $|i-j|$. Consequently, the Gaussian kernel, which is a stationary kernel (i.e., a kernel that only depends on the Euclidean distance between pairs of input points), also produces a circulant kernel matrix over the dataset. The kernel matrix is thus fully characterized by its frequency content in the Fourier domain (a circulant matrix becomes diagonal in the Fourier basis). This affords us a geometric interpretation of the generalization behavior of the kernel in the spectral (Fourier) domain.
We derive in appendix A the following equation for the prediction error $\varepsilon_{s}$ of the GP on a missing point in one orbit, which we refer to in the following as the spectral error:
$\varepsilon_{s}=\frac{\lambda^{-1}_{N}}{\langle\lambda^{-1}\rangle},$ (7)
[…]
Geometric interpretation.
@@
The formula of the spectral error lends itself to a simple geometric interpretation (Fig. 2C,D). Consider the spectrum obtained by applying the DFT to any one row of the kernel matrix. [^2] Such row is composed entirely of real entries. As a consequence, the spectrum is symmetric; we thus only show the positive half of the spectrum in the figure depictions, ranging from 0 to $N-1$. We plot the inverse spectrum $\lambda^{-1}$. In this picture, the ratio in Eq. 7 corresponds to a ratio between areas, respectively those of a rectangle with height $\lambda_{N}^{-1}$ for the numerator, and the area under the $\lambda^{-1}$ curve for the denominator.
We now turn to two case studies, in which we present further geometrical interpretations of the spectral formula (Eq. 7), which will prove useful to understand the generalization behavior of deep networks on symmetric datasets.
Effect of varying $\Delta$.
@@
We study the impact of the distance between orbits, $\Delta$, on the prediction error (Fig. 3A). We consider the dataset $\mathcal{D}_{o}$ and vary $\Delta$, as we keep the number of points in one orbit, $N$, and the kernel’s length scale, $L$, fixed. We compute the prediction error by solving the GP regression for varying $\Delta$, and compare it with our formula for the spectral error (Fig. 3B). As expected by their mathematical equivalence, the two agree across the investigated range of values, and decrease as a function of $\Delta$. This decrease aligns with the following intuition: pulling the classes apart makes prediction over the missing point by a local kernel easier.
@@
In Fig. 3C, we show the geometric interpretation of our spectral error. Increasing $\Delta$, by increasing the power of the highest frequency of the kernel matrix, skews the ratio in favor of the denominator in Eq. 7 (denoted by the red area). For $\Delta=0$, the points are interleaved in the $yz$ plane, and a kernel that would successfully generalize on the missing point would need to be dominated by its highest frequency (oscillating) component. This is not the case for the RBF kernel, whose spectrum is instead dominated by the low frequency components due to its local nature. As $\Delta$ increases, however, the kernel’s spectrum effectively gains power in the highest frequency component, thus “aligning” its properties to the requirements imposed by the regression problem.
@@
In summary, an increase in the distance between orbits $\Delta$ leads to an increase in the power of the highest frequency term $\lambda_{N}$ of the kernel matrix, which leads in turn to better kernel generalization on the missing point. This insight will prove useful in interpreting the generalization capabilities of a deep network on a high dimensional dataset possessing symmetries.
Effect of varying orbit density.
@@
We next study the impact of the number of angles, $N$, in an orbit on the prediction error (Fig. 3C). We compute the prediction error by solving GP regression for varying $N$, and compare it with our formula for the spectral error (Fig. 3D). As expected by their mathematical equivalence, the two agree across the investigated range of values, and decrease as a function of $N$. This decrease aligns with the following intuition: increasing the point density of an orbit makes prediction over a missing point easier.
@@
In Fig. 3F, we show the geometric interpretation of the spectral error. Increasing $N$ skews the ratio in favor of the denominator of Eq. 7 (denoted by the red area). The DFT projects an $n$ dimensional signal over $n$ discrete frequencies. For a local kernel, such as the RBF, which has most of its power slotted onto its first few low frequencies, sampling more points and then taking the DFT effectively means expanding the number of frequencies with low power, leading to the corresponding spectrum elements becoming vanishingly small. This in turn causes the average of the inverse spectrum to diverge, making the denominator of Eq. 7 grow. We note, however, that the presence of a nonzero $\Delta$ implies that the highest frequency terms in general, and $\lambda_{N}$ in particular, remain large, leading to the “explosion in the middle” of the inverse spectrum, which keeps the numerator in Eq. 7 small.
@@
In summary, we see that increasing the number of angles along an orbit $N$ leads to an increase in the average inverse spectrum $\langle\lambda^{-1}\rangle$ (denominator of Eq. 7), and thus to a decrease in generalization error on the missing point. Again, this insight will prove useful in interpreting the generalization capabilities of a deep network on a high dimensional dataset possessing symmetries.
3.3 Multi-layer perceptrons (MLP) on rotated MNIST
Figure 4: Analysis of the prediction behavior of a MLP on pairs of orbits from rotated-MNIST. A: We take samples of two different MNIST digits and generate their rotation orbits. The task is to predict the label value of a leave-out point in one of the two orbits. We use angle steps of 45 degrees, amounting to $N_{rot}=8$. We show the points in a reduced-dimensional space obtained by performing PCA on the dataset. B: Scatter plot of the NTK error computed in the standard way (exact) against our spectral error (Eq. 7). Each dot corresponds to a different dataset, obtained by randomly picking pairs of digits of different classes. The line of equality is superimposed as a reference. The color coding corresponds to the number of classification errors incurred by the symmetrized NTK regression, obtained by excluding a point from class A, then from class B, and counting the number of classification errors (0, 1 or 2), understood as a disagreement in sign between NTK prediction and label of the missing point. C: Comparison between the spectral and exact NTK error across different values of $\lambda^{-1}_{N}$ and $\langle\lambda^{-1}\rangle$. D: Comparison between the values of $\lambda_{N}$ and $\Delta^{2}$. The former is the highest frequency component of the neural kernel matrix (inverse of the numerator in Eq. 7), while the latter is the distance in pixel space between the averages (centroids) of the two orbits. E: Comparison between the values of $\langle\lambda^{-1}\rangle$ and $1/\rho$. The former is the average frequency component of the neural kernel matrix (denominator in Eq. 7), while $\rho$ is obtained by taking the distance of each sample in both orbits from their respective centroid, and averaging them. This quantity can be interpreted as a sort of “average density” of the two involved orbits.

### v2

@@ 3 Illustration of a Symmetry Learning Problem
Figure 1: Common deep network architectures fail to extrapolate symmetries from a partially observed version of rotated-MNIST. A: A conceptual sketch of the learning task. For samples in the leave-out class (digit “5” in this example), the upright pose is not included in the training set (empty circles). For samples in the remaining classes, all poses are included in the training set (full circles). A model that can generalize rotations should classify the missing upright samples correctly. B: The accuracies of the three models tested respectively on the training set, in-test set (normal i.i.d. test set) and out-test set (i.e., left-out pose of the leave-out class). Chance levels (10%) are also reported as a baseline. Error shades represent 95% confidence intervals computed over all tested leave-out classes (n = 10). C: Out-test accuracy for the three models as a function of the number of angles in the rotation orbit. Error shades as in B.
@@
To illustrate the problem of symmetry learning, we train a range of networks with different architectures (MLP, ConvNet, ViT-S (Dosovitskiy et al., 2021)) on partially observed versions of rotated-MNIST (details of the architectures in App. D).
@@
We start by constructing the rotated-MNIST dataset: for all samples in MNIST, we generate $N$ rotated samples, where the rotation angles are multiples of $2\pi/N$. We choose a digit class, which we refer to as the leave-out class (Fig. 1A). We define the “training” and the “in-test” splits as a 90/10 random split of the whole rotated MNIST dataset, except the upright samples of the leave-out class. These samples, instead, constitute what we call the “out-test” split. [^2] Note that predicting the labels of the out-test split is an out-of-distribution (OOD) generalization task. In this work, we study whether networks are able to perform this OOD task in virtue of the symmetric nature of the dataset.
@@
We find that increasing the number of angles leads to better out-test performance across models (Fig. 1C). Successful generalization is achieved when the number of angles sampled is sufficiently large.
In the following, we propose a theory to understand the generalization abilities of deep networks on problems involving learning from partially observed symmetries. Fully developed, our theory gives a precise account of the generalization behavior of conventionally trained networks on rotated-MNIST (Section 4.5 and Fig. 6).
4 A Neural Kernel Theory of Symmetry Learning
Essentially, the argument of this paper relies on the two following observations: (1) in the Neural Tangent Kernel (NTK) limit, training deep networks is analog to performing kernel regression; (2) kernel regression has a greatly simplified expression on datasets presenting symmetries, allowing a simple geometric interpretation of the factors leading to generalization (or lack thereof). We provide in Appendix A a brief and intuitive description of NTK. We present below the main theoretical result of the paper, which consists in the simplified expression for kernel regression on a symmetric dataset.
4.1 General Definitions
We start by recalling basic definitions relating to kernel methods.
Definition 1 (Kernel Function)
A kernel function $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ takes two inputs $x,x^{\prime}\in\mathcal{X}$ and returns a scalar value representing their similarity. It can be defined as the inner product between the images of the inputs under a certain feature map $\varphi:\mathcal{X}\to\mathcal{H}$:
$k(x,x^{\prime})=\langle\varphi(x),\varphi(x^{\prime})\rangle$
Where:
  • $\varphi(x)$ is a mapping from the original input space $\mathcal{X}$ to a higher-dimensional feature space $\mathcal{H}$ (possibly infinite-dimensional).
[…]
Knowing the kernel function allows one to skip explicit computation of the mapping $\varphi(x)$ and directly obtain the inner product.
Definition 2 (Gram Matrix)
Given a dataset $\{x_{1},x_{2},\dots,x_{N}\}$ with $N$ data points, and a kernel function $k(x,x^{\prime})$, the Gram matrix $K$ is defined as the matrix of pairwise kernel evaluations between all pairs of data points:
$K_{ij}=k(x_{i},x_{j})$
This matrix captures the pairwise relationships between all data points in the feature space induced by the kernel.
Definition 3 (Circulant Gram Matrix)
A Gram matrix $K$ is said to have a circulant structure if each row of the matrix is a cyclic shift of the previous row. Specifically, the matrix $K$ is circulant if its entries satisfy the following condition:
$K_{ij}=c_{(i-j)\mod N}$
where $N$ is the number of data points (or the size of the matrix), and $c_{k}$ represents the entries of a vector $\mathbf{c}=\{c_{0},c_{1},\dots,c_{N-1}\}$, which is periodic with period $N$. This means the matrix $K$ can be written in the form:
$K=\begin{bmatrix}c_{0}&c_{1}&c_{2}&\dots&c_{N-1}\\ c_{N-1}&c_{0}&c_{1}&\dots&c_{N-2}\\ c_{N-2}&c_{N-1}&c_{0}&\dots&c_{N-3}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ c_{1}&c_{2}&c_{3}&\dots&c_{0}\end{bmatrix}$
In other words, the matrix entries are determined by a single vector $\mathbf{c}$, and each row is a shifted version of this vector, where the shifts are cyclic (i.e., they wrap around the matrix).
4.2 Spectral Error
When the Gram matrix is circulant over a dataset made of two interleaved classes, we can derive a simplified formula for kernel regression.
Proposition 4
Consider a kernel function and its Gram matrix on a dataset. The dataset has an even number of points $2N$, where the data points are from two classes ordered in an interleaved manner with labels $+1$ and $-1$. Assume that the Gram matrix is circulant on this ordered dataset. Recalling that a circulant matrix is diagonalized by the Discrete Fourier Transform (DFT), kernel regression on a missing point from this dataset will result in the following error, referred to as the spectral error:
$\varepsilon_{s}=\frac{\lambda_{N}^{-1}}{\langle\lambda^{-1}\rangle},$ (1)
where $\lambda_{N}$ denotes the eigenvalue of the largest frequency of the Gram matrix and $\langle\lambda^{-1}\rangle$ denotes the average of the inverse eigenvalues over all frequencies (from $-N$ to $N$).
The result above, proven in App. B, will be leveraged throughout this study to gain insights into the geometrical quantities that matter for generalization of deep networks on symmetric datasets. Essentially, we will show that—on symmetric datasets—various kernels (RBF, MLP, CNN) lead to circulant (or approximately circulant) Gram matrices, allowing to use the spectral error formula (Eq. 1) to interpret the geometric factors on which generalization depends. Note that the spectral error can also be extended to multiple missing points (see App. B.1) but it leads to a less interpretable formula, so we do not consider this case.
The spectral error formula can also be generalized to any finite group, including non-abelian groups, see Theorem 15 in App. C. The resulting formula is necessarily more complex, requiring notions of representation theory and non-commutative harmonic analysis (generalized Fourier transform). For the sake of simplicity, we limit our treatment to the cyclic group in what follows.
@@ 4.3 A Simple Case Study: Gaussian Kernel Regression on a Circular Dataset
We first study the behavior of the Gaussian (i.e., RBF) kernel on a simple symmetric dataset in $\mathbb{R}^{3}$ (Fig. 2). We will see later that this case study captures all the relevant phenomenology for understanding the behavior of deep networks on datasets presenting symmetries.
Circular dataset in $\mathbb{R}^{3}$.
Consider the following dataset of $2N$ points (see Fig. 2A for an illustration):
$\mathcal{D}\equiv\{(g^{i}.x^{A},+1)\}_{i=0}^{N-1}\cup\{(g^{i}.x^{B},-1)\}_{i=0}^{N-1}\subset\mathbb{R}^{n}\times\mathbb{R},$
@@
where $g$ is a representation of the generator of the cyclic group of order $N$, acting on $\mathbb{R}^{3}$ by usual matrix multiplication. This dataset is composed of two orbits of the same group, obtained for two different seed samples, $x^{A}$ and $x^{B}$, which are labeled +1 and -1 respectively. We order the dataset so that points labeled +1 are interleaved with those labeled -1, while preserving individual orbit ordering. [^3] While this ordering does not affect the results of kernel regression on a missing point (kernel methods are order-independent), it will allow us to diagonalize the Gram matrix via the Fourier transform, instead of a permutation of it, making the spectral formula directly interpretable in terms of the eigenvalues associated to each Fourier component. For notational convenience, we denote $x_{i}^{\bullet}=g^{i}.x^{\bullet}$. The ordered dataset can be written:
$\mathcal{D}_{o}$$\equiv\{(g^{0}.x^{A},+1),(g^{0}.x^{B},-1),(g^{1}.x^{A},+1),\cdots(g^{N-1}.x^{B},-1)\}$
$\equiv\{(x_{0}^{A},+1),(x_{0}^{B},-1),(x_{1}^{A},+1),\cdots(x_{N-1}^{B},-1)\}.$
We consider the following $\mathbb{R}^{3}$ representation of the generator $g$:
$g\equiv\begin{bmatrix}1&0&0\\ 0&\cos\theta&-\sin\theta\\ 0&\sin\theta&\cos\theta\end{bmatrix},$
where we define $\theta=2\pi/N$. We consider the seed points
$x^{A}\equiv\frac{\Delta}{2}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}+\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},\quad x^{B}\equiv-\frac{\Delta}{2}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}+\begin{bmatrix}0\\ \cos(\theta/2)\\ \sin(\theta/2)\end{bmatrix},$ (2)
@@
over which we make $g$ act by usual matrix multiplication. Successive applications of $g$ onto the seed points generate the dataset. Note that the second seed point $x^{B}$ is chosen such that the two orbits are geometrically interleaved (as depicted in fig 2A), which means that points of orbit $A$ are equidistant from their two nearest neighbors in orbit $B$ and vice-versa. This geometric interleaving is a necessary condition for the Gaussian kernel Gram matrix to be circulant (see below).
@@
Figure 2: Case study: Analysis of the generalization behavior of the Gaussian kernel on a circular-symmetric dataset. A: A circular dataset is made of two sets of interleaved points in $\mathbb{R}^{3}$ belonging to two classes (denoted in red and blue). One point (in black) is left out during training, and we regress on it. B: The pairwise distance matrix between the points is circulant (above), leading to a circulant kernel matrix (below) obtained by applying the Gaussian (RBF) kernel function to it elementwise. We select the $N$-th row from it, highlighted by the red rectangle; due to the circularity of the matrix, this comes without loss of information. C: A plot of the selected row of the kernel matrix. The values of the kernel alternate between two limiting curves, representing respectively the “same label” and the “opposite label” kernel values. This alternation, due to the separation between classes, is akin to the presence of a high frequency component in the kernel function itself, were it computed over the classes perfectly interleaved ($\Delta$ = 0). D: The prediction error of a Gaussian kernel on a missing point of a circular-symmetric dataset is a simple function of its spectrum (Eq. 1). The reciprocal (inverse) of the positive half of the DFT of the selected row of the kernel matrix is shown. The grey area corresponds to the numerator of the spectral error, while the red area corresponds to the denominator. The ratio of the two corresponds to the prediction error incurred by kernel regression using the chosen kernel. The presence of the aforementioned high frequency component in the kernel is manifested in the low value of $\lambda_{N}^{-1}$, leading to small error (i.e., increased class separability).
@@
Figure 3: Further geometric interpretations of the Gaussian kernel generalization behavior on a circular-symmetric dataset. A: We progressively increase the separation $\Delta$ between two circular classes of points (respectively in blue and red). The leave-out point used for testing is in black. B: The prediction error (in black) of the RBF kernel on the leave-out point decreases as a function of $\Delta$, as predicted by our spectral formula (in red). given in Eq. 1. C: Inverse spectra of the kernel matrix for different $\Delta$. The grey (respectively, red) area corresponds to the numerator (resp., denominator) of Eq. 1. As $\Delta$ increases, the ratio between grey and red areas (i.e., the spectral error) progressively decreases, as a consequence of the last frequencies getting larger. D, E, F: Same as A,B,C where instead of increasing $\Delta$ we increase the number of points forming each class. The ratio between grey and red areas progressively decreases with number of points, as a consequence of the middle inverse frequencies progressively diverging.
Gaussian (RBF) kernel regression on a circular dataset.
@@
We consider the following regression problem: remove any one point from the dataset $\mathcal{D}_{o}$, and estimate the value of $y$ at the missing point. We solve this regression problem using Gaussian process (GP) kernel regression. We select the Gaussian or Radial Basis Function (RBF) kernel,
$k_{RBF}(x_{i},x_{j})=\exp(-L^{2}\lVert x_{i}-x_{j}\rVert_{2}^{2}),$
where $L$ denotes the kernel’s length scale.
Lemma 5
The Gram matrix of a stationary kernel (a kernel that only depends on the Euclidean distance between pairs of input points) over a dataset made of two geometrically interleaved cyclic orbits (as defined above) is circulant.
Proof  First consider that the pairwise distance between points $x_{i}$ and $x_{j}$ in $\mathcal{D}_{o}$ is only function of their absolute index difference $|i-j|$. Since a stationary kernel only depends on the Euclidean distance between pairs of input points, it follows that a stationary kernel produces a circulant Gram matrix over $\mathcal{D}_{o}$.
The RBF kernel is a stationary kernel. The Gram matrix of the RBF kernel being circulant over $\mathcal{D}_{o}$, Proposition 4 applies, and prediction at the missing point is given by the spectral error (Eq. 1). This affords us a geometric interpretation of the generalization behavior of the kernel in the spectral (Fourier) domain.
Geometric interpretation
@@
The formula of the spectral error lends itself to a simple geometric interpretation (Fig. 2). Consider the spectrum obtained by applying the DFT to any one row of the kernel matrix. [^4] Such row is composed entirely of real entries. As a consequence, the spectrum is symmetric; we thus only show the positive half of the spectrum in the figure depictions, ranging from 0 to $N-1$. We plot the inverse spectrum $\lambda^{-1}$. In this picture, the ratio in Eq. 1 corresponds to a ratio between areas, respectively those of a rectangle with height $\lambda_{N}^{-1}$ for the numerator, and the area under the $\lambda^{-1}$ curve for the denominator.
@@
Let us now study the impact of the distance between orbits, $\Delta$, on the prediction error (Fig. 3A-C). We vary $\Delta$ while keeping the number of points in one orbit, $N$, and the kernel’s length scale, $L$, fixed. We compute the prediction error by solving kernel regression for varying $\Delta$, and compare it with our formula for the spectral error (Fig. 3B). As expected by their mathematical equivalence, the two agree across the investigated range of values, and decrease as a function of $\Delta$. This decrease aligns with the following intuition: pulling the classes apart makes prediction over the missing point by a local kernel easier.
@@
In Fig. 3C, we show the geometric interpretation of the spectral error. Increasing $\Delta$, by increasing the power of the highest frequency of the kernel matrix, skews the ratio in favor of the denominator in Eq. 1 (denoted by the red area). For $\Delta=0$, the points are interleaved in the $yz$ plane, and a kernel that would successfully generalize on the missing point would need to be dominated by its highest frequency (oscillating) component. This is not the case for the RBF kernel, whose spectrum is instead dominated by the low frequency components due to its local nature. As $\Delta$ increases, however, the kernel’s spectrum effectively gains power in the highest frequency component, thus “aligning” its properties to the requirements imposed by the regression problem.
@@
In summary, an increase in the distance between orbits $\Delta$ leads to an increase in the power of the highest frequency term $\lambda_{N}$ of the kernel matrix, which leads in turn to better kernel generalization on the missing point.
@@
We now turn to study the impact of the number of angles in an orbit $N$ on the prediction error (Fig. 3D-F). We compute the prediction error by solving kernel regression for varying $N$, and compare it with our formula for the spectral error (Fig. 3E). As expected by their mathematical equivalence, the two agree across the investigated range of values, and decrease as a function of $N$. This decrease aligns with the following intuition: increasing the point density of an orbit makes prediction over a missing point easier.
@@
In Fig. 3F, we show the geometric interpretation of the spectral error. Increasing $N$ skews the ratio in favor of the denominator of Eq. 1 (denoted by the red area). The DFT projects an $n$ dimensional signal over $n$ discrete frequencies. For a local kernel, such as the RBF, which has most of its power slotted onto its first few low frequencies, sampling more points and then taking the DFT effectively means expanding the number of frequencies with low power, leading to the corresponding spectrum elements becoming vanishingly small. This in turn causes the average of the inverse spectrum to diverge, making the denominator of Eq. 1 grow. [^5] We note, however, that the presence of a nonzero $\Delta$ implies that the highest frequency terms in general, and $\lambda_{N}$ in particular, remain large, leading to the “explosion in the middle” of the inverse spectrum, which keeps the numerator in Eq. 1 small.
@@
In summary, we see that increasing the number of angles along an orbit $N$ leads to an increase in the average inverse spectrum $\langle\lambda^{-1}\rangle$ (denominator of Eq. 1), and thus to a decrease in generalization error on the missing point.
These geometric insights will prove useful in interpreting the generalization capabilities of deep networks on high dimensional datasets possessing symmetries.
4.4 Multi-layer Perceptrons on Rotated-MNIST